# Spin drag mechanism of giant thermal magnetoresistance

### v2

@@ Abstract
In this work, we introduce a different drag-induced thermal effect. We show that at charge neutrality, spin polarization of the electrons strongly affects the thermal conductivity. The salient feature of heat transport at charge neutrality is that it can proceed via the hydrodynamic flow of the neutral electron-hole plasma. We thus focus our consideration on the hydrodynamic regime, which attracted significant attention in recent years, see reviews Lucas and Fong (2018); Levchenko and Schmalian (2020); Narozhny (2022) and references therein. In a pristine system, the thermal conductivity becomes infinite, and thus very sensitive to disorder and other perturbations. In realistic systems, the thermal conductivity is limited by the disorder-induced friction, which can be made sufficiently weak in high mobility systems. Significant hydrodynamic enhancement of thermal conductivity in graphene systems has been reported in Ref. Crossno et al. (2016). Continuing progress in nanofabrication enables fabrication of samples with an even greater mobility, making the thermal conductivity of the system extremely sensitive to other perturbations. Here, we show that thermal transport properties of high mobility systems become very sensitive to spin polarization of the liquid, leading to strong thermal magnetoresistivity. The microscopic mechanism of this phenomenon in the hydrodynamic regime involves spin diffusion (or spin drag). The physical reason for this effect can be understood by observing that since thermal conductivity is measured at zero spin-current, the convective part of the spin current must be compensated by spin-diffusion relative to the liquid. This dramatically increases dissipation at nonzero spin polarization, resulting in giant thermal magnetoresistivity. Our predictions may enable probing spin drag in electron-hole plasma in graphene by thermal measurements Waissman et al. (2022).
@@
with $\eta$ and $\zeta$ being, respectively, shear (first) and bulk (second) viscosity of the electron liquid. The force density in the right hand side of Eq. (4) comes from the local gradients of pressure in the fluid $P$. To express it in this form we used the thermodynamic relation $\bm{\nabla}P=s\bm{\nabla}T+\varsigma\bm{\nabla}\mu_{\sigma}=\vec{x}^{\mathbb{T}}\vec{\bm{X}}$ and the column vector notations of Eq. (2). The superscript $\mathbb{T}$ denotes transposition. The remaining term in Eq. (4) describes the generic disorder-induced friction characterized by the friction coefficient $k$. Working under the assumption of smooth disorder, namely weak disorder potential with the long correlation radius $\xi$, the coefficient of friction $k$ can be expressed in terms of the local density variations $\delta n(\bm{r})$ induced by disorder potential and the intrinsic conductivity $\sigma$ as follows Li et al. (2020)
@@
In closing we note that our consideration focused on the bulk contribution to thermal spin drag magnetotransport where momentum relaxation is driven by the disorder potential and the hydrodynamic flow velocity is uniform. In devices whose dimensions are smaller or comparable to the Gurzhi length there will be additional contribution to the thermal resistance, which is determined by the viscous flow near sample boundaries. An extension of the present theory to the devices with Hall bar and Corbino geometry will be presented elsewhere.
This research project was financially supported by the National Science Foundation Grant No. DMR-2203411 and H. I. Romnes Faculty Fellowship provided by the University of Wisconsin-Madison Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation (A. L.). A. V. A. acknowledges support from the National Science Foundation through the MRSEC Grant No. DMR-1719797, the Thouless Institute for Quantum Matter (TIQM), and the College of Arts and Sciences at the University of Washington.

### v3

@@ Abstract
In this work, we introduce a different drag-induced thermal effect. We show that near charge neutrality, spin polarization of the electrons strongly affects the thermal conductivity. The salient feature of heat transport at charge neutrality is that it can proceed via the hydrodynamic flow of the neutral electron-hole plasma. We thus focus our consideration on the hydrodynamic regime, which attracted significant attention in recent years, see reviews Lucas and Fong (2018); Levchenko and Schmalian (2020); Narozhny (2022) and references therein. In a pristine system, the thermal conductivity becomes infinite, and thus very sensitive to disorder and other perturbations. In realistic systems, the thermal conductivity is limited by the disorder-induced friction, which can be made sufficiently weak in high mobility systems. Significant hydrodynamic enhancement of thermal conductivity in graphene systems has been reported in Ref. Crossno et al. (2016). Continuing progress in nanofabrication enables fabrication of samples with an even greater mobility, making the thermal conductivity of the system extremely sensitive to other perturbations. Here, we show that thermal transport properties of high mobility systems become very sensitive to spin polarization of the liquid, leading to strong thermal magnetoresistivity. The microscopic mechanism of this phenomenon in the hydrodynamic regime involves spin diffusion (or spin drag). The physical reason for this effect can be understood by observing that since thermal conductivity is measured at zero spin-current, the convective part of the spin current must be compensated by spin-diffusion relative to the liquid. This dramatically increases dissipation at nonzero spin polarization, resulting in giant thermal magnetoresistivity. Our predictions may enable probing spin drag in electron-hole plasma in graphene by thermal measurements Waissman et al. (2022); Talanov et al. (2024).
@@
with $\eta$ and $\zeta$ being, respectively, shear (first) and bulk (second) viscosity of the electron liquid. The force density in the right hand side of Eq. (4) comes from the local gradients of pressure in the fluid $P$. To express it in this form we used the thermodynamic relation $\bm{\nabla}P=s\bm{\nabla}T+\varsigma\bm{\nabla}\mu_{\sigma}=\vec{x}^{\mathbb{T}}\vec{\bm{X}}$ and the column vector notations of Eq. (2). The superscript $\mathbb{T}$ denotes transposition. The remaining term in Eq. (4) describes the generic disorder-induced friction characterized by the friction coefficient $k$. For weak disorder whose correlation radius $\xi$ exceeds the electron inelastic mean free path $l_{\text{ee}}$, the coefficient of friction $k$ can be expressed in terms of the local density variations $\delta n(\bm{r})$ induced by disorder potential and the intrinsic conductivity $\sigma$ as follows Li et al. (2020)
@@
To evaluate magneto-thermal conductivity induced by spin drag away from charge neutrality it is important to realize that the right-hand-side of Eq. (8) must be evaluated at vanishing spin and electric current. Although the latter condition is no longer automatically satisfied at nonvanishing particle density $n$, for $n\ll s$ the result can be obtained by a straightforward extension of the above consideration. This requires two main modifications. First, the Coulomb force $-en\bm{E}$, where the electromotive force (EMF) $e\bm{E}$ is the gradient of the local electrochemical potential, must be added to the right-hand-side of the force balance condition, Eq. (4). Second, the constitutive relations in Eq. (1) have to be augmented to include the following expression for the electric current density, $\bm{j}_{e}=en\bm{u}+\sigma\bm{E}-\frac{\gamma}{T}\bm{\nabla}T$, where $\gamma$ is the intrinsic thermoelectric coefficient. At $n\ll s$ the thermoelectric contribution to $\bm{j}_{e}$ may be neglected because $\gamma/T\sim n/s\ll 1$. In this case, the condition of vanishing current gives $\bm{E}=-(en/\sigma)\bm{u}$. Inserting this expression into the force-balance condition gives a force term $(en)^{2}\bm{u}/\sigma$, which amounts to the following redefinitoin of the effective friction coefficient in the previous expressions,
$k\to k(n)=k+\frac{e^{2}}{\sigma}n^{2}.$ (15)
Thus, the dependence of the relative thermal magnetoresistance in Eq. (14) on $n$ has the form of the Lorentzian with a width $\sim\sqrt{\langle\delta n^{2}\rangle}$, as follows from Eq. (6). The limit of charge neutrality corresponds to $n\ll\delta n$. In modern monolayer graphene devices density inhomogeneity ranges between $\sqrt{\langle\delta n^{2}\rangle}\sim(5\div 10)\times 10^{9}$ cm-2 Crossno et al. (2016); Talanov et al. (2024); Xin et al. (2023). This is still safely compatible with the condition $\{n,\sqrt{\langle\delta n^{2}\rangle}\}<s$ at $T\sim 100$ K in monolayer graphene where hydrodynamic behavior of Dirac plasma is expected to emerge.
In closing, we note that our consideration focused on the bulk contribution to thermal spin drag magnetotransport where momentum relaxation is driven by the disorder potential and the hydrodynamic flow velocity is uniform. In devices whose dimensions are smaller or comparable to the Gurzhi length there will be additional contribution to the thermal resistance, which is determined by the viscous flow near sample boundaries. An extension of the present theory to the devices with Hall bar and Corbino geometry will be presented elsewhere.
We thank Philip Kim for useful discussion. This research project was financially supported by the National Science Foundation Grant No. DMR-2203411 and H. I. Romnes Faculty Fellowship provided by the University of Wisconsin-Madison Office of the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation (A. L.). A. V. A. acknowledges support from the National Science Foundation through the MRSEC Grant No. DMR-1719797, the Thouless Institute for Quantum Matter (TIQM), and the College of Arts and Sciences at the University of Washington.